OpenStudy (anonymous):

Calculate the slope of the line between the points (9,-50) and (2, -8)

OpenStudy (anonymous):

-8 - (-50)/2-9. Integer rules apply here, -8 - (-50) = 42. The subtraction sign and the negative sign in front of the 50 cancel out each other, making it 50 - 8. 2- 9 = -7. So 42/-7. We can't have negative signs on the bottom of the fraction. So multiply the entire thing by -1/-1 making it -42/7. Reduce which makes it -6. The slope is -6.

OpenStudy (anonymous):

You are so good...I am amazed!

OpenStudy (anonymous):

Calculate the slope of the line between the points (9,-62) and (5,-30)

OpenStudy (anonymous):

Is the answer 8?

OpenStudy (anonymous):

-8 I think?

OpenStudy (anonymous):

-30 - (-62)/5-9. -30 - (-62), same as before the subtraction sign and the negative sign in front of the 62 cancel each other out making it 62 - 30 = 32. 5 - 9 = -4, so the fraction looks like this so far: 32/-4. multiply by -1/-1 making it -32/4. The answer is -8, close though - you just forgot your negative sign :)

OpenStudy (anonymous):

(2,8) (1,-7) I'm doing this one and I'm coming up with 1, but that answer isn't there??

OpenStudy (anonymous):

-7 - 8 = -15 because integer rules make the subtraction sign into a plus sign and the 8 becomes negative. 1 - 2 = -1. -15/-1 *-1/-1 = 15/1 which becomes 15. Answer is 15, do you know what you did wrong??

OpenStudy (anonymous):

Now i do ! Thanks

OpenStudy (anonymous):

(7,-6)(-3,-2)

OpenStudy (anonymous):

-5/2 ?